# SePer: Measure retrieval utility through the lens of semantic perplexity reduction

### v4

@@ 1 Introduction
Retrieval plays a crucial role in satisfying information needs across various interactive systems. With the rapid advancement of Large Language Models (LLMs), retrieval has become deeply interwoven with generation processes Lewis et al. (2020); Guu et al. (2020). This integration not only enhances the accuracy and faithfulness of generated content Chen et al. (2017) but also enables handling more complex applications such as multi-hop reasoning Trivedi et al. (2022a), visual information seeking Hu et al. (2024), and task completion Yao et al. (2023); Zhang et al. (2023).
@@ 2 Related Works
Evaluation of Information Retrieval. The effectiveness of retrieval is often evaluated from two aspects: Direct content evaluation, which scrutinizes the relevance of the retrieved content itself, and response-based evaluation, which gauges the quality of the responses to reflect on the effectiveness of in-the-middle modules Salemi & Zamani (2024). However, each method suffers from several shortcomings individually. Direct context evaluation treats retrieval as an independent module, which cannot reflect the impact on the receiver model. Response-based methods can be further divided into two categories: model-based and reference-based methods. Model-based methods require a knowledgable model, such as human and GPT-4 Liu et al. (2023), to evaluate whether the LLM output is a desired response to a given query. Reference-based methods require a reference answer and evaluate the LLM outputs by their matching score to the reference, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and BERTScore Zhang et al. (2020). However, these methods mainly reflect lexical information and cannot capture the nuanced relationship in semantic meaning. More recent works also leverage LLM judges, such as ARES Saad-Falcon et al. (2024), Prometheus Kim et al. (2024), and Auto-J Li et al. (2024) to assess the distance between reference and output. These LLM-based methods are often more accurate due to the language understanding ability and comprehensive world knowledge of LLMs. However, they are often slow and expensive.
@@ 4.1.1 Validity of semantic-based answer scoring
Datasets. We use EVOUNA Wang et al. (2024), a Question Answering (QA) benchmark to evaluate the reliability of QA evaluators. Based on Natural Questions (NQ) Kwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017), EVOUNA augmented the datasets with LLM-generated responses and asked humans to annotate whether a response is semantically equivalent to the golden answer.
@@ 4.1.2 Validity of $\Delta$SePer on quantifying retrieval utility
Baselines. We choose various methods that can be used to estimate the LLM’s knowledge and use their difference to measure retrieval utility. Lexical-matching-based methods include EM, ROUGE, and BLEU, which measure the response correctness score through matching text spans. BERTScore matches predicted answers and ground truth through embedding similarity. Another category of baselines is uncertainty measures, such as perplexity, entropy, and semantic entropy. Unlike the matching-based method, these uncertainty measures do not require golden answers in computation. While they are not directly defined on the correctness dimension, we include them because recent literature also shows that in calibrated LLMs, uncertainty is correlated with knowledge capabilities Farquhar et al. (2024); Cheng et al. (2024). We also included LLM-judges similar to Table 1.
@@ A.1 Dataset Statistics
This section presents the datasets used in our experiments, detailing their names, sizes, sources, and key characteristics.
Dataset Size Source Key Characteristics
@@
MS MARCO Bajaj et al. (2016) 101,093 Bing Search engine queries, web passages
PopQA Mallen et al. (2023) 14,267 Wikipedia Focuses on popular culture questions
TriviaQA Joshi et al. (2017) 11,313 Wikipedia & Web Trivia questions with evidence documents
SQuAD Rajpurkar et al. (2016) 10,570 Wikipedia Standard reading comprehension
NQ Kwiatkowski et al. (2019) 3,610 Wikipedia Real user queries, diverse topics
BoolQ Clark et al. (2019) 3,270 Wikipedia Yes/No questions, requires inference
AmbigQA Min et al. (2020) 2,002 Wikipedia Question ambiguity with multiple answers
Fermi Kalyan et al. (2021) 1,000 Wikipedia Estimation-based reasoning
@@
HotpotQA Yang et al. (2018) 7,405 Wikipedia Requires multi-hop reasoning
2WikiMultihopQA Ho et al. (2020) 12,576 Wikipedia Cross-document multi-hop reasoning
MuSiQue Trivedi et al. (2022b) 2,417 Wikipedia Multi-hop reasoning, complex questions
Bamboogle Press et al. (2023) 125 Wikipedia Compositional reasoning challenging for Google
@@
FEVER Thorne et al. (2018) 10,444 Wikipedia Fact verification
@@
MMLU Hendrycks et al. (2021) 14,042 N/A Multiple-choice, general knowledge
@@
WikiASP Hayashi et al. (2021) 37,368 Wikipedia Open-domain summarization
Table 3: Dataset Details
@@ C.1 Benchmarking Retriever Quality and Utility
We evaluate six dense retrievers: $\textit{AAR}_{\textit{ANCE}}$ Yu et al. (2023), $\textit{AAR}_{\textit{Contriever}}$ Yu et al. (2023), BGE Xiao et al. (2024), Contriever Izacard et al. (2022), DPR Karpukhin et al. (2020), and E5 Wang et al. (2022), alongside the classical sparse retriever BM25 Robertson et al. (1995). The results of the quality and utility benchmarks are presented in Tables 10 and 11, respectively.
@@ D.2 Rerankers
    $\text{score}(q,d_{i})=f\left(\mathbf{h}_{[\text{CLS}]}\right),$ (16)
    where $\mathbf{h}_{[\text{CLS}]}$ represents the contextual representation of the [CLS] token. Alternatively, some architectures may use pooling methods (e.g., mean or max pooling) overall token representations or token-level interactions to derive the relevance score.

### v5

@@ 1 Introduction
Retrieval plays a crucial role in satisfying information needs across various interactive systems. With the rapid advancement of Large Language Models (LLMs) and various downstream applications Xu et al. (2025); Ye et al. (2025); Guo et al. (2025), retrieval has become deeply interwoven with generation processes Lewis et al. (2020); Guu et al. (2020). This integration not only enhances the accuracy and faithfulness of generated content Chen et al. (2017) but also enables handling more complex applications such as multi-hop reasoning Trivedi et al. (2022a), visual information seeking Hu et al. (2024), and task completion Yao et al. (2023); Zhang et al. (2023).
@@ 2 Related Works
Evaluation of Information Retrieval. The effectiveness of retrieval is often evaluated from two aspects: Direct content evaluation, which scrutinizes the relevance of the retrieved content itself, and response-based evaluation, which gauges the quality of the responses to reflect on the effectiveness of in-the-middle modules Salemi & Zamani (2024). However, each method suffers from several shortcomings individually. Direct context evaluation treats retrieval as an independent module, which cannot reflect the impact on the receiver model. Response-based methods can be further divided into two categories: model-based and reference-based methods. Model-based methods require knowledgable models, such as human and GPT-4 Liu et al. (2023), to evaluate whether the LLM output is a desired response to a given query. Reference-based methods require a reference answer and evaluate the LLM outputs by their matching score to the reference, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and BERTScore Zhang et al. (2020). However, these methods mainly reflect lexical information and cannot capture the nuanced relationship in semantic meaning. More recent works also leverage LLM judges, such as ARES Saad-Falcon et al. (2024), Prometheus Kim et al. (2024), and Auto-J Li et al. (2024) to assess the distance between reference and output. These LLM-based methods are often more accurate due to the language understanding and LLMs’ comprehensive world knowledge. However, they are often slow and expensive.
@@ 4.1.1 Validity of semantic-based answer scoring
Datasets. We use EVOUNA Wang et al. (2024), a Question Answering (QA) benchmark to evaluate QA evaluators’ reliability. Based on Natural Questions (NQKwiatkowski et al. (2019) and TriviaQA Joshi et al. (2017), EVOUNA augmented the datasets with LLM-generated responses and asked humans to annotate whether a response is semantically equivalent to the golden answer.
@@ 4.1.2 Validity of $\Delta$SePer on quantifying retrieval utility
Baselines. We choose various methods that can be used to estimate the LLM’s knowledge and use their difference to measure retrieval utility. Lexical-matching-based methods include EM, ROUGE, and BLEU, which measure the response correctness score through matching text spans. BERTScore matches predicted answers and ground truth through embedding similarity. Another category of baselines is uncertainty measures, such as perplexity, entropy, and semantic entropy. Unlike the matching-based method, these uncertainty measures do not require golden answers in computation. While they are not directly defined on the correctness dimension, we include them as recent literature also shows that uncertainty is correlated with knowledge capabilities Farquhar et al. (2024); Cheng et al. (2024) in calibrated LLMs. We also included LLM-judges similar to Table 1.
@@ A.1 Dataset Statistics
This section presents the datasets used in our experiments, detailing their names, sizes, sources, key characteristics, and availability, which are categorized by task type (Single QA, Multi-hop QA, Fact Verification, Multiple-choice QA, and Summarization). The table below also includes links to the original papers and dataset repositories.
Dataset Size Source Key Characteristics Availability
@@
MS Marco Bajaj et al. (2016) 101,093 Bing Search engine queries, web passages [Paper] [Dataset]
PopQA Mallen et al. (2023) 14,267 Wikipedia Focuses on popular culture questions [Paper] [Dataset]
TriviaQA Joshi et al. (2017) 11,313 Wikipedia & Web Trivia questions with evidence documents [Paper] [Dataset]
SQuAD Rajpurkar et al. (2016) 10,570 Wikipedia Standard reading comprehension [Paper] [Dataset]
NQ Kwiatkowski et al. (2019) 3,610 Wikipedia Real user queries, diverse topics [Paper] [Dataset]
BoolQ Clark et al. (2019) 3,270 Wikipedia Yes/No questions, requires inference [Paper] [Dataset]
AmbigQA Min et al. (2020) 2,002 Wikipedia Question ambiguity with multiple answers [Paper] [Dataset]
Fermi Kalyan et al. (2021) 1,000 Wikipedia Estimation-based reasoning [Paper] [Dataset]
@@
HotpotQA Yang et al. (2018) 7,405 Wikipedia Requires multi-hop reasoning [Paper] [Dataset]
2WikiMultihopQA Ho et al. (2020) 12,576 Wikipedia Cross-document multi-hop reasoning [Paper] [Dataset]
MuSiQue Trivedi et al. (2022b) 2,417 Wikipedia Multi-hop reasoning, complex questions [Paper] [Dataset]
Bamboogle Press et al. (2023) 125 Wikipedia Compositional reasoning challenging for Google [Paper] [Dataset]
@@
FEVER Thorne et al. (2018) 10,444 Wikipedia Fact verification [Paper] [Dataset]
@@
MMLU Hendrycks et al. (2021) 14,042 N/A Multiple-choice, general knowledge [Paper] [Dataset]
@@
WikiASP Hayashi et al. (2021) 37,368 Wikipedia Open-domain summarization [Paper] [Dataset]
Table 3: Summary of Datasets Used in Our Experiments
@@ C.1 Benchmarking Retriever Quality and Utility
We evaluate six dense retrievers: $\texttt{AAR}_{\texttt{ANCE}}$ Yu et al. (2023), $\texttt{AAR}_{\texttt{Contriever}}$ Yu et al. (2023), BGE Xiao et al. (2024), Contriever Izacard et al. (2022), DPR Karpukhin et al. (2020), and E5 Wang et al. (2022), alongside the classical sparse retriever BM25 Robertson et al. (1995). The results of the quality and utility benchmarks are presented in Tables 10 and 11, respectively.
@@ D.2 Rerankers
    $\text{score}(q,d_{i})=f\left(\mathbf{h}_{\texttt{[CLS]}}\right),$ (16)
    where $\mathbf{h}_{\texttt{[CLS]}}$ represents the contextual representation of the [CLS] token. Alternatively, some architectures may use pooling methods (e.g., mean or max pooling) over all token representations or token-level interactions to derive the relevance score.